what are the x and y intercepts of the equation 2x-2y =3
Answer by fractalier(6550)

To find the x and y intercepts of the equation 2x - 2y = 3, we set x=0 to find the y-intercept...so we get
2(0) - 2y = 3
-2y = 3
y = -3/2
To find the x-intercept, we let y=0...
2x - 2(0) = 3
2x = 3
x = 3/2